# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### v2

@@ 1 Introduction
Despite their success, the neural networks can occasionally lead the agent to make suboptimal decisions that an adversarial opponent may exploit. Wang et al. (2023) have illustrated how policy networks trained via AlphaZero-style algorithms in Go are susceptible to exploitation, as evidenced by winning against the trained policy network 97% of the time. Similarly, Perolat et al. (2022) report that their trained agents display evident gameplay errors, necessitating the application of multiple heuristics to rectify these errors. However, there are principled ways of using search to avoid these mistakes, as evidenced by AlphaZero (Silver et al., 2018), DeepStack (Moravčík et al., 2017), and Student of Games (Schmid et al., 2023).
Algorithms that use policy-gradient methods in imperfect information games like Regularized Nash Dynamics (RNaD) (Perolat et al., 2022, 2021) rely on directly applying the neural network’s policy during gameplay. Adding search to these methods could improve the strategy and mitigate some mistakes. However, this enhancement comes at the expense of turning model-free algorithms into model-dependent ones during gameplay.
@@
In this paper, we introduce the algorithm Multi-agent Search with Policy Transformations (SePoT), which builds upon the concept of multi-valued states (Brown et al., 2018). SePoT trains a policy transformation critic network that estimates the expected value of policies derived from the policy network. This critic is trained alongside a policy network with a policy-gradient technique without requiring any additional trajectories beyond those sampled by the original policy-gradient algorithm. The critic network is trained with an off-policy V-trace that estimates expected values against a multitude of potential opponent strategies derived from the policy represented by the policy network. We show that this critic provides sufficient information to perform a safe search (Šustr et al., 2021) in imperfect information games.
Furthermore, our approach enables safe search at any point within a game, utilizing the information provided by the introduced critic network. This eliminates the need for conducting search in all decision points from the root until the current decision point, which is often intractable in games with little public knowledge. It is in strict contrast to prior methods (Brown et al., 2018; Moravčík et al., 2017; Schmid et al., 2023). This flexibility allows us to perform search only in game segments sufficiently simple for available compute power, while employing the neural network policy in other parts of the game.
Finally, we establish a connection between value functions and multi-valued states within imperfect information games that enables training previously used imperfect information value functions (Moravčík et al., 2017; Schmid et al., 2023) without the expensive search in the training time.
In our experiments, we evaluate SePoT with RNaD (Perolat et al., 2022). We show that using search in head-to-head play in imperfect-information Goofspiel and Battleships improves the performance over the network trained with RNaD. Furthermore, we demonstrate that the exploitability of policy derived from the search is the same or lower than that of the policy network from RNaD in a small version of imperfect information Goofspiel and the standard benchmark game Leduc hold’em.
@@ 2 Issues with imperfect information search
Consider a smaller version of Battleships, in which two players have three ships of sizes 3, 2, and 2 on a square grid of size 5. Players take turns shooting on an opponent’s board after all ships are secretly placed. The player that first hits all tiles occupied by the opponent’s ships wins. Suppose we are in the middle of the game, and the current boards are shown in Figure 1. Each player knows their remaining ship’s position, but does not know the opponent’s. Based on the public shooting actions up until this point, we know that there are only two possible positions for the player and the opponent. The player may have either placed the ship at A1-A2 or A2-B2. Similarly, the opponent may have placed the ship at A1-A3 or A3-C3. Possible positions are shown on each board with the dashed rectangle. There are four possible states, that is, all possible combinations between the positions across both players, as shown in Figure 0(c). The player cannot distinguish between states $w_{1}$ and $w_{2}$, neither between $w_{3},w_{4}$. Similarly, the opponent cannot distinguish between $w_{1},w_{3}$ and between $w_{2},w_{4}$. Shooting anywhere else than A1, A2, A3, B3, and C3 does not make sense for a rational player. However, training RNaD for three days results in a policy that shoots to these rational tiles with a probability less than $45\%$.
@@
Employing additional search should improve this policy, but doing so naively in imperfect information games introduces multiple issues. The first issue is that starting search only from states that are possible from the player’s point of view, which are $w_{1}$ and $w_{2}$, would search only in those parts of the game where the player placed the ship at A1-A2. This means that the player incorrectly assumes the opponent knows the ship’s position during this search. The outcome of the search would be that the game always ends in 2 turns because the opponent would shoot at the ship. Since the player requires at least three turns to win, it could never win against this informed opponent, and its strategy would be to play randomly. However, in the real gameplay scenario, the opponent lacks knowledge of the player’s ship’s location, providing the player, who is moving first, with a genuine opportunity to win. Nevertheless, this issue can be fixed by starting the search from all four possible states, which does not make the same error.
@@
The second issue is more profound. In practice, for many different reasons, the player may incorrectly assume that the opponent is slightly more likely to place his ship at position A1 instead of C3. In such a case, the player would always shoot at either tile A1, A2, or A3 since it provides more reward in expectation from its point of view. This leads to a strategy that could be exploited by an adversarial opponent who would never put the ship there in the first place. This is a well-studied problem in imperfect information games. One possible solution during search is to assume that the opponent can play any strategy in earlier parts of the game. This assumption then allows computing a more robust strategy against possible play from the game’s previous and subsequent parts (Burch et al., 2014; Moravcik et al., 2016; Brown and Sandholm, 2017).
@@
The third issue for applying search in the situation in Figure 1 are sizes of required game structures. The full end-game cannot be solved online in reasonable time since it has over $10^{9}$ terminal histories. Also, it is impossible to safely search in earlier decision points, since after the first move there are more than $10^{8}$ possible ship configurations and branching factor of $50$, which menas that even a depth-limited game with look-ahead of one move has more than $10^{10}$ histories. Using a depth-limited search with a value function at the depth limit only later in the game enables search even with much less computational resources (Kovařík et al., 2023).
@@ 3 Background
Factored-observation stochastic game (FOSG) (Kovařík et al., 2022) is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{W},p,w^{0},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O})$, where $\mathcal{N}=\{1,\dots,N\}$ is a player set, $\mathcal{W}$ is the set of states and $w^{0}$ is an initial state, $p:\mathcal{W}\rightarrow 2^{\mathcal{N}}$ is a player function, which for given state $w\in\mathcal{W}$ assigns currently acting players from $\mathcal{N}$, $\mathcal{A}=\Pi_{i\in\mathcal{N}}\mathcal{A}_{i}$ is the space of joint actions. We denote $\mathcal{A}_{i}(w)$ legal actions for player $i$ in the world state $w$. $\mathcal{T}:\mathcal{W}\times\mathcal{A}\rightarrow\mathcal{W}$ is the transition functiion, $\mathcal{R}:\mathcal{W}\times\mathcal{A}\rightarrow\mathbb{R}^{\mathcal{N}}$ is the reward function and $\mathcal{O}:\mathcal{W}\times\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{O}$ is the observation function. $\mathcal{O}$ is factored as $\mathcal{O}=(\mathcal{O}_{0},\mathcal{O}_{1},\dots\mathcal{O}_{N})$, where $\mathcal{O}_{0}$ are a public observations given to all players and $\mathcal{O}_{i}$ are a private observations of player $i$. In this work, we focus on two-player zero-sum games, where $\mathcal{N}=\{1,2,c\}$ and $\mathcal{R}_{1}(w,a)=-\mathcal{R}_{2}(w,a)\;\forall(w,a)\in\mathcal{W}\times\mathcal{A}$. $c$ denotes a chance player who has publicly announced a non-changeable policy throughout the game. Moreover, we study only games with perfect recall, where neither player forgets any information.
@@
Trajectory $\tau=w^{0}a^{0}w^{1}a^{1}\dots w^{t}\in(\mathcal{W}\mathcal{A})^{*}\mathcal{W}$ for which $a^{l}\in\mathcal{A}(w^{l})$ and $w^{l+1}=\mathcal{T}(w^{l},a^{l})$ is a finite sequence of states and actions in the game. Trajectory utility for player $i$ is a cumulative reward $u_{i}(\tau)=\sum_{t=0}^{l-1}\mathcal{R}_{i}(w^{t},a^{t})$. We use $\mathcal{H}$ to denote histories, which are trajectories that start in the initial state $w^{0}$. $\mathcal{Z}=\{z\in\mathcal{H}|z\text{ is terminal}\}$ are histories in which neither player has an action to play, effectively ending the game. $h\sqsubseteq h^{\prime}$ means that $h^{\prime}$ extends $h$ and $h^{0}$ is an initial history, that corresponds to the initial state $w^{0}$, before any player played any action. Each history ends with some world state $w^{l}$. Therefore, we denote $\mathcal{R}_{i}(h,a)=\mathcal{R}_{i}(w^{l},a)$ as a reward given to player $i$, if in history $h\in\mathcal{H}$ the players play joint action $a$. Since player $i$ does not observe the state of the game directly, but only observes its actions $\mathcal{A}_{i}$, private observations $\mathcal{O}_{i}$ and public observations $\mathcal{O}_{0}$, it cannot distinguish between multiple different histories. We say that these indistinguishable histories are in the same infoset $s_{i}\in\mathcal{S}_{i}$. If two histories $h,h^{\prime}\in\mathcal{H}$ belong to the same infoset $s_{i}$, then player $i$ has to have the same actions in these histories and $s_{i}(h)$ denotes an infoset of player $i$ in history $h$. We will use $\mathcal{A}_{i}(s_{i})$ to denote these available actions in infoset $s_{i}$. Furthermore, we will use a notion of public-states $s_{0}\in\mathcal{S}_{0}$, which may be viewed as an information set for an external player which does not act in the game. Two histories belong to the same public state if they have the same sequence of public observations from the initial state. Each public state is made from multiple information sets for a given player $i$. We will use $\mathcal{S}_{i}(s_{0})$ to denote all infosets that are in public state $s_{0}$ and $\mathcal{H}(s)$ to denote all histories in an infoset $s$.
@@
Expected utility, if all players play according to strategy profile $\pi$ from history $h$ onward is defined as $u^{\pi}_{i}(h)=\mathbb{E}_{\tau\sim\pi|h}u_{i}(\tau)$. $P^{\pi}(h)=\prod_{h^{\prime}aw\sqsubseteq h}\prod_{j\in\mathcal{N}}\pi_{j}(s_{j}(h^{\prime}),a_{j})$ is a probability of reaching history $h$ if all players play according to strategy profile $\pi$. This may be separated into $P^{\pi}(h)=P^{\pi}_{i}(h)P^{\pi}_{-i}(h)$, where $P^{\pi}_{i}(h)$ is only a players $i$ contribution to reaching history $h$, similarly $P^{\pi}_{-i}(h)$ is a contribution of all the other players except $i$, which is called counterfactual reach probability. Counterfactual values for infoset $s_{i}$ is a sum of expected utilities of each history $h$ within the same infoset, weighted by the counterfactual reach probability of this history $v_{i}^{\pi}(s_{i})=\frac{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)u_{i}^{\pi}(h)}{\sum_{s_{i}(h)=s_{i}}P^{\pi}_{-i}(h)}$.
@@ 3.1 Depth-limited solving
Resolving some subgame to improve blueprint policy $\pi$ naively would involve fixing policy $\pi$ for all players in previous parts of the game and using single $w^{A}$, where players act with corresponding ranges. However, this may result in a much more exploitable strategy (Burch et al., 2014) than the original blueprint $\pi$. Multiple techniques were developed to deal with this issue in two-player zero-sum games (Burch et al., 2014; Moravcik et al., 2016; Brown and Sandholm, 2017; Brown et al., 2020). In this work we will focus on the one proposed in (Burch et al., 2014). When resolving from the perspective of player $i$, it constructs a slightly modified subgame called a gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$, where $v^{\pi}_{o}$ are counterfactual values for each infoset $s_{o}\in\mathcal{S}_{o}(s_{0})$ of opponent $o$ in a current public state $s_{0}$, when players play according to policy $\pi$. This subgame has the artificial state $w^{A}$ as an initial state, but instead of using full ranges $r$, it uses only ranges $r_{i}$ and $r_{c}$. Directly after this artificial state, $|\mathcal{H}(s_{0})|$ opponent’s decision nodes are separated into multiple infosets based on the information player has in a given history. The opponent $o$ has a choice to either terminate the game and receive the counterfactual value $v^{\pi}_{o}(s_{o})$ or continue the game as in the original subgame.
Suppose that after the depth limit, the optimal value function is used and that the counterfactual values $v^{\pi^{{NE}}}_{o}(s_{o})$ used in the gadget game are computed for Nash equilibrium policy $\pi^{{NE}}$. Then the solution of gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$ is part of a Nash equilibrium in the entire game (Burch et al., 2014; Kovařík et al., 2023).
In this work, we focus on an alternative to the value function called multi-valued states (Brown et al., 2018), which, after the depth limit, gives the opponent the choice to fix its strategy until the end of the game. Then, the game terminates with the reward corresponding to the reward if player $i$ follows some blueprint policy $\pi_{i}$, while the opponent follows the fixed strategy. Authors of (Kovařík et al., 2023) pointed out that multi-valued states could be viewed as a form of a value function. We further discuss this in Section 4.2

### v3

@@ 1 Introduction
Despite their success, the neural networks can occasionally lead the agent to make suboptimal decisions that an adversarial opponent may exploit. AlphaZero-style algorithms in Go are susceptible to exploitation, as evidenced by winning against the trained policy network 97% of the time Wang et al. (2023). Similarly, policy-gradient algorithms display evident gameplay errors, necessitating the application of multiple heuristics to rectify these errors Perolat et al. (2022). However, there are principled ways of using search to avoid these mistakes, as evidenced by AlphaZero Silver et al. (2018), DeepStack Moravčík et al. (2017), and Student of Games Schmid et al. (2023).
Algorithms that use policy-gradient methods in imperfect information games like Regularized Nash Dynamics (RNaD) Perolat et al. (2022, 2021) rely on directly applying the neural network’s policy during gameplay. Adding search to these methods could improve the strategy and mitigate some mistakes. However, this enhancement comes at the expense of turning model-free algorithms into model-dependent ones during gameplay.
@@
In this paper, we introduce the algorithm Multi-Agent Search With Policy Transformations (SePoT), which builds upon the concept of multi-valued states Brown et al. (2018). SePoT trains a policy transformation critic network that estimates the expected value of policies derived from the policy network. This critic is trained alongside a policy network with a policy-gradient technique without requiring any additional trajectories beyond those sampled by the original policy-gradient algorithm. The critic network is trained with an off-policy V-trace that estimates expected values against a multitude of potential opponent strategies derived from the policy represented by the policy network. We show that this critic provides sufficient information to perform a safe search Šustr et al. (2021) in imperfect information games.
Furthermore, our approach enables safe search at any point within a game, utilizing the information provided by the introduced critic network. This eliminates the need for conducting search in all decision points from the root until the current decision point, which is often intractable in games with little public knowledge, but at the cost of requiring a domain-specific function that reconstructs current game segment. It is in strict contrast to prior methods Brown et al. (2018); Moravčík et al. (2017); Schmid et al. (2023). This flexibility allows us to perform search only in game segments sufficiently simple for available compute power, while employing the neural network policy in other parts of the game.
Finally, we establish a connection between value functions and multi-valued states within imperfect information games that enables training previously used imperfect information value functions Moravčík et al. (2017); Schmid et al. (2023) without the expensive search in the training time.
In our experiments, we evaluate SePoT with RNaD Perolat et al. (2022). We show that using search in head-to-head play in imperfect-information Goofspiel and Battleships improves the performance over the network trained with RNaD. Furthermore, we demonstrate that the exploitability of policy derived from the search is the same or lower than that of the policy network from RNaD in a small version of imperfect information Goofspiel and the standard benchmark game Leduc hold’em.
@@ 2 Issues With Imperfect Information Search
Consider a smaller version of Battleships, in which two players have three ships of sizes 3, 2, and 2 on a square grid of size 5. Players take turns shooting on a tile of the opponent’s board after all ships are secretly placed. The player that first hits all tiles occupied by the opponent’s ships wins. Suppose we are in the middle of the game, and the current boards are shown in Figure 1. Each player knows their remaining ship’s position, but does not know the opponent’s. Based on the public shooting actions up until this point, we know that there are only two possible positions for the player and the opponent. The player may have either placed the ship at A1-A2 or A2-B2. Similarly, the opponent may have placed the ship at A1-A3 or A3-C3. Possible positions are shown on each board with the dashed rectangle. There are four possible states, that is, all possible combinations between the positions across both players, as shown in Figure 1(c). The first player cannot distinguish between states $w_{1}$ and $w_{2}$, neither between $w_{3},w_{4}$. Similarly, the opponent cannot distinguish between $w_{1},w_{3}$ and between $w_{2},w_{4}$. Shooting anywhere else than A1, A2, A3, B3, and C3 does not make sense for a rational player. However, training RNaD for three days results in a policy that shoots to these tiles with a probability less than $45\%$.
@@
Employing additional search should improve this policy, but doing so naively in imperfect information games introduces multiple issues. The first issue is that starting search only from states that are possible from the player’s point of view, which are $w_{1}$ and $w_{2}$, would search only in those parts of the game where the player placed the ship at A1-A2. This means that the player incorrectly assumes the opponent knows the ship’s position during this search. The outcome of the search would be that the game always ends in 2 turns because the opponent would shoot at the ship. Since the player requires at least three turns to win, it could never win against this informed opponent, and its strategy would be to play randomly. However, in the real gameplay scenario, the opponent lacks knowledge of the player’s ship’s location, providing the player, who is moving first, with a genuine opportunity to win. Nevertheless, this issue can be fixed by starting the search from all four possible states.
@@
The second issue is more profound. In practice, for many different reasons, the player may incorrectly assume that the opponent is slightly more likely to place his ship at position A1 instead of C3. In such a case, the player would always shoot at either tile A1, A2, or A3 since it provides more reward in expectation from its point of view. This leads to a strategy that could be exploited by an adversarial opponent who would never put the ship there in the first place. This is a well-studied problem in imperfect information games. One possible solution during search is to assume that the opponent can play any strategy in earlier parts of the game. This assumption then allows computing a more robust strategy against possible play from the game’s previous and subsequent parts Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017).
@@
The third issue for applying search in the situation in Figure 1 are sizes of required game structures. The full end-game cannot be solved online in reasonable time since it has over $10^{9}$ terminal histories. Also, it is impossible to safely search in earlier decision points, since after the first move there are more than $10^{8}$ possible ship configurations and branching factor of $50$, which means that even a depth-limited game with look-ahead of one move has more than $10^{10}$ histories. Using a depth-limited search with a value function at the depth limit only later in the game enables search even with much less computational resources Kovařík et al. (2023).
@@ 3 Background
Factored-observation stochastic game (FOSG) Kovařík et al. (2022) is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{W},p,w^{0},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O})$, where $\mathcal{N}=\{1,\dots,N\}$ is a player set, $\mathcal{W}$ is the set of states and $w^{0}$ is an initial state, $p:\mathcal{W}\rightarrow 2^{\mathcal{N}}$ is a player function, which for given state $w\in\mathcal{W}$ assigns currently acting players from $\mathcal{N}$, $\mathcal{A}=\Pi_{i\in\mathcal{N}}\mathcal{A}_{i}$ is the space of joint actions. We denote $\mathcal{A}_{i}(w)$ legal actions for player $i$ in the world state $w$. $\mathcal{T}:\mathcal{W}\times\mathcal{A}\rightarrow\mathcal{W}$ is the transition function, $\mathcal{R}:\mathcal{W}\times\mathcal{A}\rightarrow\mathbb{R}^{\mathcal{N}}$ is the reward function and $\mathcal{O}:\mathcal{W}\times\mathcal{A}\times\mathcal{W}\rightarrow\mathbb{O}$ is the observation function. $\mathcal{O}$ is factored as $\mathcal{O}=(\mathcal{O}_{0},\mathcal{O}_{1},\dots\mathcal{O}_{N})$, where $\mathcal{O}_{0}$ are a public observations given to all players and $\mathcal{O}_{i}$ are a private observations of player $i$. In this work, we focus on two-player zero-sum games, where $\mathcal{N}=\{1,2,c\}$ and $\mathcal{R}_{1}(w,a)=-\mathcal{R}_{2}(w,a)\;\forall(w,a)\in\mathcal{W}\times\mathcal{A}$. $c$ denotes a chance player who has publicly announced a non-changeable policy throughout the game. Moreover, we study only games with perfect recall, where neither player forgets any information.
@@
Trajectory $\tau=w^{0}a^{0}w^{1}a^{1}\dots w^{t}\in(\mathcal{W}\mathcal{A})^{*}\mathcal{W}$ for which $a^{l}\in\mathcal{A}(w^{l})$ and $w^{l+1}=\mathcal{T}(w^{l},a^{l})$ is a finite sequence of states and actions in the game. Trajectory utility for player $i$ is a cumulative reward $u_{i}(\tau)=\sum_{t=0}^{l-1}\mathcal{R}_{i}(w^{t},a^{t})$. We use $\mathcal{H}$ to denote histories, which are trajectories that start in the initial state $w^{0}$. $\mathcal{Z}=\{z\in\mathcal{H}|z\text{ is terminal}\}$ are histories in which neither player has an action to play, effectively ending the game. $h\sqsubseteq h^{\prime}$ means that $h^{\prime}$ extends $h$ and $h^{0}$ is an initial history, that corresponds to the initial state $w^{0}$, before any player played any action. Each history ends with some world state $w^{l}$. Therefore, we denote $\mathcal{R}_{i}(h,a)=\mathcal{R}_{i}(w^{l},a)$ as a reward given to player $i$, if in history $h\in\mathcal{H}$ the players play joint action $a$. Since player $i$ does not observe the state of the game directly, but only observes its actions $\mathcal{A}_{i}$, private observations $\mathcal{O}_{i}$ and public observations $\mathcal{O}_{0}$, it cannot distinguish between multiple different histories. We say that these indistinguishable histories are in the same infoset $s_{i}\in\mathcal{S}_{i}$. If two histories $h,h^{\prime}\in\mathcal{H}$ belong to the same infoset $s_{i}$, then player $i$ has to have the same actions in these histories and $s_{i}(h)$ denotes an infoset of player $i$ in history $h$. We will use $\mathcal{A}_{i}(s_{i})$ to denote these available actions in infoset $s_{i}$. Furthermore, we will use a notion of public-states $s_{0}\in\mathcal{S}_{0}$, which may be viewed as an information set for an external player which does not act in the game. Two histories belong to the same public state if and only if they have the same sequence of public observations from the initial state. Each public state is made from multiple information sets for a given player $i$. We will use $\mathcal{S}_{i}(s_{0})$ to denote all infosets that are in public state $s_{0}$ and $\mathcal{H}(s)$ to denote all histories in an infoset $s$.
@@
Expected utility, if all players play according to strategy profile $\pi$ from history $h$ onward is defined as $u^{\pi}_{i}(h)=\mathbb{E}_{\tau\sim\pi|h}u_{i}(\tau)$. $P^{\pi}(h)=\prod_{h^{\prime}aw\sqsubseteq h}\prod_{j\in\mathcal{N}}\pi_{j}(s_{j}(h^{\prime}),a_{j})$ is a probability of reaching history $h$ if all players play according to strategy profile $\pi$. This may be separated into $P^{\pi}(h)=P^{\pi}_{i}(h)P^{\pi}_{-i}(h)$, where $P^{\pi}_{i}(h)$ is only a players $i$ contribution to reaching history $h$, similarly $P^{\pi}_{-i}(h)$ is a contribution of all the other players except $i$, which is called counterfactual reach probability. Counterfactual values for infoset $s_{i}$ is a sum of expected utilities of each history $h$ within the same infoset, weighted by the counterfactual reach probability of this history $v_{i}^{\pi}(s_{i})=\frac{\sum_{h\in\mathcal{H}(s_{i})}P^{\pi}_{-i}(h)u_{i}^{\pi}(h)}{\sum_{h\in\mathcal{H}(s_{i})}P^{\pi}_{-i}(h)}$.
@@ 3.1 Depth-Limited Solving
Resolving some subgame to improve blueprint policy $\pi$ naively would involve fixing policy $\pi$ for all players in previous parts of the game and using single $w^{A}$, where players act with corresponding ranges. However, this may result in a much more exploitable strategy Burch et al. (2014) than the original blueprint $\pi$. Multiple techniques were developed to deal with this issue in two-player zero-sum games Burch et al. (2014); Moravcik et al. (2016); Brown and Sandholm (2017); Brown et al. (2020). In this work we will focus on the one proposed in Burch et al. (2014). When resolving from the perspective of player $i$, it constructs a slightly modified subgame called a gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$, where $v^{\pi}_{o}$ are counterfactual values for each infoset $s_{o}\in\mathcal{S}_{o}(s_{0})$ of opponent $o$ in a current public state $s_{0}$, when players play according to policy $\pi$. This subgame has the artificial state $w^{A}$ as an initial state, but instead of using full ranges $r$, it uses only ranges $r_{i}$ and $r_{c}$. Directly after this artificial state, $|\mathcal{H}(s_{0})|$ opponent’s decision nodes are separated into multiple infosets based on the information player has in a given history. The opponent $o$ has a choice to either terminate the game and receive the counterfactual value $v^{\pi}_{o}(s_{o})$ or continue the game as in the original subgame.
Suppose that after the depth limit, the optimal value function is used and that the counterfactual values $v^{\pi^{{NE}}}_{o}(s_{o})$ used in the gadget game are computed for Nash equilibrium policy $\pi^{{NE}}$. Then the solution of gadget game $\mathcal{G}^{G}_{i}(\beta,v^{\pi}_{o})$ is part of a Nash equilibrium in the entire game Burch et al. (2014); Kovařík et al. (2023).
In this work, we focus on an alternative to the value function called multi-valued states Brown et al. (2018), which, after the depth limit, gives the opponent the choice to fix its strategy until the end of the game. Then, the game terminates with the reward corresponding to the reward if player $i$ follows some blueprint policy $\pi_{i}$, while the opponent follows the fixed strategy. Authors of Kovařík et al. (2023) pointed out that multi-valued states could be viewed as a form of a value function. We further discuss this in Section 4.2